# Quantum refraction effects in pulsar radio emission

### v1

@@ Abstract
Highly magnetized neutron stars exhibit the vacuum non-linear electrodynamics effects, which can be well described using the one-loop effective action for quantum electrodynamics. In this context, we study the propagation and polarization of pulsar radio emission, based on the post-Maxwellian Lagrangian from the Heisenberg-Euler-Schwinger action. Given the refractive index obtained from this Lagrangian, we determine the leading-order corrections to both the propagation and polarization vectors due to quantum refraction via perturbation analysis. In addition, the effects on the orthogonality between the propagation and polarization vectors and the Faraday rotation angle, all due to quantum refraction are investigated. Furthermore, from the dual refractive index and the associated polarization modes, we discuss quantum birefringence, with the optical phenomenology analogous to its classical counterpart.
@@
pulsars – magnetic fields – radio emission – non-linear electrodynamics – post-Maxwellian Lagrangian model – quantum refraction – quantum birefringence
pubyear: 2024pagerange: Quantum refraction effects in pulsar radio emissionA
@@ 1 Introduction
In this paper, we study the propagation and polarization of a photon in the dipole magnetic field background of a pulsar model, based on the post-Maxwellian (PM) Lagrangian; it is, in the weak field limit, the generic form of non-Maxwellian Lagrangian for the non-linear vacuum (Sorokin (2022)) and has been used to test vacuum polarization effects in the PVLAS (Polarizzazione del Vuoto con LAser) project (Ejlli et al. (2020)). The Heisenberg-Euler-Schwinger (HES) Lagrangian of strong-field QED (Heisenberg & Euler (1936); Schwinger (1951)) is also well approximated by the PM Lagrangian below the critical field strength. Here, we develop the recent analysis of vacuum birefringence in Kim & Kim (2023) further for a practical model of pulsar curvature radio emission, where the magnetic field is defined for an oblique rotator with an inclination angle and the electric field is discarded. We consider the dual refractive index and the associated polarization vectors of a probe photon (Kim & Kim (2022)) to investigate the photon propagation in the strong magnetic field background of this pulsar model. Then the leading-order corrections to both the propagation and polarization vectors due to quantum refraction are determined via perturbation analysis. Our study provides a novel, complementary approach to and an elaboration of other similar studies in Heyl & Shaviv (2000, 2002); Heyl et al. (2003); Caiazzo & Heyl (2018); Heyl & Caiazzo (2018); Caiazzo (2019); Caiazzo & Heyl (2021); Caiazzo et al. (2022).
@@
The paper is organized as follows. In Section 2 we work out the deflection of a light ray from pulsar emission due to quantum refraction; the leading-order corrections to the propagation vector and then to the trajectory of the light ray are determined. In Section 3 we look into the dual refractive index and the associated polarization modes of the light ray under the effect of quantum refraction; the leading-order corrections to the polarization vectors for Case I (Section 3.1) and Case II (Section 3.2) are determined. In addition, the effects on the orthogonality between the propagation and polarization vectors and the Faraday rotation angle, due to quantum refraction are investigated. Furthermore, in regard to the optical phenomenology from the dual refractive index and the associated polarization modes, we discuss quantum birefringence for this pulsar emission. Then finally, we conclude the paper with discussions on other similar studies and follow-up studies.
@@ 2 Deflection of a light ray due to quantum refraction
The expression of $\mathbf{\hat{n}}$ for $n\neq\mathrm{const.}$ can be applied to a mechanism of how the light ray is deflected, for example, due to the quantum refraction effect in pulsar radio emission, as will be described below. In the presence of the effect, the refractive index $n$ is given by a function of the position $\mathbf{r}$, at which the light ray crosses a local magnetic field line in a pulsar magnetosphere; otherwise, it would simply be a constant.
According to the PM Lagrangian, the refractive index $n$ is given by (Kim & Kim (2022))
@@
where $\eta_{1}$ and $\eta_{2}$ are parameters defined via $\eta_{1}/4=\eta_{2}/7=e^{4}/\left(360\pi^{2}m_{\mathrm{e}}^{4}\right)$, and $B$ is the local magnetic field strength at a point in a pulsar magnetosphere, and $\vartheta$ denotes the angle between the light ray trajectory and the local magnetic field line (see Fig. 1). Here we have named Case I and Case II for two different values of the refractive index attributed to the same magnetosphere, depending on the propagation and polarization of the light ray associated with them. However, for the rest of this Section, we focus on Case I as there is little difference in the propagation of the light ray between the two cases. Then in Section 3 we look into the polarization of the light ray for both the cases and discuss quantum birefringence in relation to it.
One can expand the refractive index $n$ for Case I in equation (4), having $\eta_{1}$, $\eta_{2}\sim 10^{-31}\,\mathrm{g}^{-1}\,\mathrm{cm\,s}^{2}$ and $B<B_{\mathrm{c}}$ (critical magnetic field)$\,\sim 10^{13}\,\mathrm{G}$, and thus $\eta_{1}B^{2}$, $\eta_{2}B^{2}\ll 1$. Then it can be approximated as
@@
Figure 1: A cross-sectional view of a pulsar magnetosphere with the dipole magnetic field lines around a neutron star. The vertical dashed line (black) and the inclined solid line (red) represent the rotation axis and the magnetic axis, respectively. $\alpha$ between these axes denotes the inclination angle. The field lines in curved spacetime (blue) are distinguished from those in flat spacetime (green); in this paper, only the latter are relevant to our analysis as we rule out the effect of gravitation due to the neutron star mass. The scale of the unity in this graph is equivalent to the neutron star radius $\sim 10^{6}\,\mathrm{cm}$. The red dashed curve represents the trajectory of the light ray traced by $\mathbf{\hat{n}}_{\left[0\right]}$ as projected onto the $xz$-plane. (Credit: Kim & Trippe (2021), reproduced with modifications.)
@@
In our radio emission model, a light ray of curvature radiation is emitted from the dipole magnetic field:
@@
$\mathbf{\hat{n}}_{\left[0\right]}=\beta\mathbf{\hat{B}}+\mathbf{\Omega}\times\mathbf{r,}$ (8)
@@
$\hat{n}_{y\left[0\right]}\approx\frac{\Omega}{c}\left[\left(\hat{n}_{x\left[0\right]}+\frac{\sin\alpha}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}\right)s+x_{\mathrm{o}}\right]+\mathcal{O}\left(\phi^{2},\left(\Omega r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega r_{\mathrm{o}}/c\right)\right),$ (12)
@@
$y\approx\int\hat{n}_{y\left[0\right]}\,\mathrm{d}s+y_{\mathrm{o}}=\frac{\Omega}{c}\left[\frac{1}{2}\left(\hat{n}_{x\left[0\right]}+\frac{\sin\alpha}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}\right)s+x_{\mathrm{o}}\right]s+y_{\mathrm{o}},$ (15)
@@
$\hat{n}_{y}=\hat{n}_{y\left[0\right]}\approx\frac{\Omega r_{\mathrm{o}}}{c}\left[\left(\hat{n}_{x\left[0\right]}+\frac{\sin\alpha}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}\right)\lambda+\sin\theta_{\mathrm{o}}\right],$ (26)
@@
$Y\equiv\int_{0}^{s}\delta\hat{n}_{y}\,\mathrm{d}s=r_{\mathrm{o}}\int_{0}^{\lambda}\hat{n}_{y\left[0\right]}\,\mathrm{d}\lambda\approx\frac{\Omega r_{\mathrm{o}}^{2}}{c}\left[\frac{1}{2}\left(\hat{n}_{x\left[0\right]}+\frac{\sin\alpha}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}\right)\lambda+\sin\theta_{\mathrm{o}}\right]\lambda,$ (29)
@@
$\left.\hskip 33.0pt+8\cos\left(\theta_{\mathrm{o}}-\alpha\right)\left(141\cos^{4}\left(\theta_{\mathrm{o}}-\alpha\right)+86\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+13\right)\!\!\right]$
@@
For example, with $r_{\mathrm{o}}=2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\Omega=2\pi\times 10^{2}\,\mathrm{Hz}$ and $\eta_{2}B_{\mathrm{o}}^{2}\sim 10^{4}$ (fairly exaggerated for the purpose of visualization, $10^{9}$ times as large as an actual value $\sim 10^{-5}$), we obtain a trajectory curve of our light ray $\left(X/r_{\mathrm{o}},Y/r_{\mathrm{o}},Z/r_{\mathrm{o}}\right)$ for $0\leq\lambda\leq 10$, as shown in Fig. 2 below. Note, in particular, that the trajectory is deflected from a straight line as viewed in the $xz$-plane (due to the quantum refraction effect), and at the same time that it follows a parabolic path in another plane perpendicular to the $xz$-plane (due to the rotational effect of the pulsar magnetosphere); therefore, the light ray follows a three-dimensional twisted curve. In Appendix A we provide a detailed discussion of the properties of this curve in reference to the Frenet–Serret formulas (Spivak (1999)).
Figure 2: Trajectories of the light ray $\left(X/r_{\mathrm{o}},Y/r_{\mathrm{\ o}},Z/r_{\mathrm{o}}\right)$ plotted against $0\leq\lambda\leq 10$; the red solid curve represents the total trajectory (classical trajectory $+$ quantum correction), while the blue and green dashed curves represent the classical trajectory and the quantum correction, respectively. For the purpose of visualization, the effect is fairly exaggerated here by $\eta_{2}B_{\mathrm{o}}^{2}\sim 10^{4}$, which is $10^{9}$ times as large as an actual value $\sim 10^{-5}$.

### v2

# Quantum refraction effects in pulsar emission

@@ Abstract
Highly magnetized neutron stars exhibit the vacuum non-linear electrodynamics effects, which can be well described using the one-loop effective action for quantum electrodynamics. In this context, we study the propagation and polarization of pulsar radiation, based on the post-Maxwellian Lagrangian from the Heisenberg-Euler-Schwinger action. Given the refractive index obtained from this Lagrangian, we determine the leading-order corrections to both the propagation and polarization vectors due to quantum refraction via perturbation analysis. In addition, the effects on the orthogonality between the propagation and polarization vectors and the Faraday rotation angle, all due to quantum refraction are investigated. Furthermore, from the dual refractive index and the associated polarization modes, we discuss quantum birefringence, with the optical phenomenology analogous to its classical counterpart.
@@
pulsars – magnetic fields – curvature radiation – non-linear electrodynamics – post-Maxwellian Lagrangian model – quantum refraction – quantum birefringence
pubyear: 2024pagerange: Quantum refraction effects in pulsar emissionA
@@ 1 Introduction
In this paper, we study the propagation and polarization of a photon in the dipole magnetic field background of a pulsar model, based on the post-Maxwellian (PM) Lagrangian; it is, in the weak field limit, the generic form of non-Maxwellian Lagrangian for the non-linear vacuum (Sorokin (2022)) and has been used to test vacuum polarization effects in the PVLAS (Polarizzazione del Vuoto con LAser) project (Ejlli et al. (2020)). The Heisenberg-Euler-Schwinger (HES) Lagrangian of strong-field QED (Heisenberg & Euler (1936); Schwinger (1951)) is also well approximated by the PM Lagrangian below the critical field strength. Here, we develop the recent analysis of vacuum birefringence in Kim & Kim (2023) further for a practical model of pulsar emission, where the magnetic field is defined for an oblique rotator with an inclination angle and the electric field is discarded. We consider the dual refractive index and the associated polarization vectors of a probe photon (Kim & Kim (2022)) to investigate the photon propagation in the strong magnetic field background of this pulsar model. Then the leading-order corrections to both the propagation and polarization vectors due to quantum refraction are determined via perturbation analysis. Our study provides a novel, complementary approach to and an elaboration of other similar studies in Heyl & Shaviv (2000, 2002); Heyl et al. (2003); Caiazzo & Heyl (2018); Heyl & Caiazzo (2018); Caiazzo (2019); Caiazzo & Heyl (2021); Caiazzo et al. (2022).
@@
The paper is organized as follows. In Section 2 we review the Lagrangian formalism for the non-linearity of the quantum vacuum due to strong electromagnetic fields; a brief account of the HES Lagrangian and the PM Lagrangian as its approximation is given. In Section 3 we work out the deflection of a light ray from pulsar emission due to quantum refraction; the leading-order corrections to the propagation vector and then to the trajectory of the light ray are determined. In Section 4 we look into the dual refractive index and the associated polarization modes of the light ray under the effect of quantum refraction; the leading-order corrections to the polarization vectors for Case I (Section 4.1) and Case II (Section 4.2) are determined. In addition, the effects on the orthogonality between the propagation and polarization vectors and the Faraday rotation angle, due to quantum refraction are investigated. Furthermore, in regard to the optical phenomenology from the dual refractive index and the associated polarization modes, we discuss quantum birefringence for this pulsar emission. Then finally, we conclude the paper with discussions on other similar studies and follow-up studies.
2 Non-linear Electrodynamics due to Strong Fields
Non-linear electrodynamic (NED) effects of the vacuum in the presence of strong electromagnetic fields have been studied employing the effective Lagrangian formalism. One of the most well-known NED models is the Heisenberg-Euler-Schwinger (HES) Lagrangian, which is obtained by adding the one-loop QED correction to the Maxwell Lagrangian, due to spin-1/2 fermions of mass $m_{\mathrm{e}}$ and charge $e$ in electromagnetic fields of arbitrary strengths (Heisenberg & Euler (1936); Schwinger (1951)): in the convention with $\hbar=c=1$,
$\mathcal{L}_{\mathrm{HES}}\left(a,b\right)=\mathcal{L}^{\left(0\right)}\left(a,b\right)+\mathcal{L}^{\left(1\right)}\left(a,b\right)=\frac{b^{2}-a^{2}}{2}+\mathcal{L}^{\left(1\right)}\left(a,b\right),$ (1)
where $\mathcal{L}^{\left(0\right)}\left(a,b\right)$ refers to the classical Maxwell Lagrangian, defined through
$a\equiv\sqrt{\sqrt{F^{2}+G^{2}}+F},~{}b\equiv\sqrt{\sqrt{F^{2}+G^{2}}-F},$ (2)
with the Lorentz- and gauge-invariant Maxwell scalar $F$ and pseudo-scalar $G$
$F\equiv\frac{1}{4}F^{\mu\nu}F_{\mu\nu}=\frac{1}{2}\left(\mathbf{B}^{2}-\mathbf{E}^{2}\right)=\frac{1}{2}\left(a^{2}-b^{2}\right),~{}G\equiv\frac{1}{4}F^{\mu\nu}F_{\mu\nu}^{\ast}=-\mathbf{E}\cdot\mathbf{B}=\mathrm{sign}\left(\pm\right)ab,$ (3)
and $\mathcal{L}^{\left(1\right)}\left(a,b\right)$ refers to the Lagrangian of one-loop correction,
$\mathcal{L}^{\left(1\right)}\left(a,b\right)=-\frac{1}{8\pi^{2}}\int_{0}^{\infty}\mathrm{d}s\,\frac{e^{-m_{\mathrm{e}}^{2}s}}{s^{3}}\left\{\left(eas\right)\coth\left(eas\right)\left(ebs\right)\cot\left(ebs\right)-\left[1+\frac{\left(eas\right)^{2}-\left(ebs\right)^{2}}{3}\right]\right\}.$ (4)
However, in the weak-field limit (below the critical field strength $B_{\mathrm{c}}$), the HES Lagrangian (1) has the leading-order contribution, the so-called post-Maxwellian (PM) Lagrangian (Euler & Kockel (1935)): [^1] In effect, the PM Lagrangian approximates the HES Lagrangian well even up to $B=0.14B_{\mathrm{c}}$, with an error less than 1%.
$\mathcal{L}_{\mathrm{PM}}\left(a,b\right)=-F+\eta_{1}F^{2}+\eta_{2}G^{2},$ (5)
where $\eta_{1}$ and $\eta_{2}$ are parameters defined via $\eta_{1}/4=\eta_{2}/7=e^{4}/\left(360\pi^{2}m_{\mathrm{e}}^{4}\right)$.
In some NED models, a parity violating term proportional to $FG$ is added to the PM Lagrangian (Ni et al. (2013)). However, for the rest of the paper, our analysis is based on the PM Lagrangian; the dual refractive index as given by equation (9) in Section 3 is derived using this (Adler (1971); Kim & Kim (2022)).
@@ 3 Deflection of a light ray due to quantum refraction
The expression of $\mathbf{\hat{n}}$ for $n\neq\mathrm{const.}$ can be applied to a mechanism of how the light ray is deflected, for example, due to the quantum refraction effect in pulsar emission, as will be described below. In the presence of the effect, the refractive index $n$ is given by a function of the position $\mathbf{r}$, at which the light ray crosses a local magnetic field line in a pulsar magnetosphere; otherwise, it would simply be a constant.
According to Kim & Kim (2022), the refractive index $n$ can be derived using the PM Lagrangian (5) as
@@
where $B$ is the local magnetic field strength at a point in a pulsar magnetosphere, and $\vartheta$ denotes the angle between the light ray trajectory and the local magnetic field line (see Fig. 1). Here we have named Case I and Case II for two different values of the refractive index attributed to the same point in the magnetosphere; the propagation and polarization of the light ray are associated with these values. Later in Sections 4.1 and 4.2, the two polarization vectors for Case I and Case II, as given by (41) and (50), respectively are set to be orthogonal to each other and to the propagation vector such that the three vectors form a classical orthonormal basis. However, for the rest of this Section, we focus on Case I as there is little difference in the propagation of the light ray between the two cases. Then in Section 4 we look into the polarization of the light ray for both the cases and discuss quantum birefringence in relation to it.
One should note that the refractive index $n$ as given by (9) has no dependence on the frequency of radiation. Consequently, in our entire analysis, the quantum refraction effects derived from this, on the propagation and polarization of a photon have no frequency dependence either, as can be checked with equations (11) and (40), respectively later. However, in order for equation (9) to be considered valid, the frequency of pulsar radiation must be significantly lower than that for excitation of the quantum vacuum ($\sim 10^{20}\,\mathrm{Hz}$; in the gamma-ray regime), which corresponds to the photon energy required to create an electron-positron pair, such that the vacuum is far from resonance. In addition, the plasma effects may be neglected if the pulsar radiation frequency is much higher than the local plasma frequency ($\sim 10^{9}\,\mathrm{Hz}$) (Petri (2016)). Therefore, our pulsar radiation can be safely assumed to cover optical to X-ray emissions ($\sim 10^{12}$ to $10^{17}\,\mathrm{Hz}$) in this work.
One can expand the refractive index $n$ for Case I in equation (9), having $\eta_{1}$, $\eta_{2}\sim 10^{-31}\,\mathrm{g}^{-1}\,\mathrm{cm\,s}^{2}$ and $B<B_{\mathrm{c}}$ (critical magnetic field)$\,\sim 10^{13}\,\mathrm{G}$, and thus $\eta_{1}B^{2}$, $\eta_{2}B^{2}\ll 1$. Then it can be approximated as (Adler (1971))
@@
Figure 1: A cross-sectional view of a pulsar magnetosphere with the dipole magnetic field lines (green) around a neutron star. The vertical dashed line (black) and the inclined solid line (red) represent the rotation axis and the magnetic axis, respectively. $\alpha$ between these axes denotes the inclination angle. The scale of the unity in this graph is equivalent to the neutron star radius $\sim 10^{6}\,\mathrm{cm}$. The red dotted line represents the trajectory curve of the light ray traced by $\mathbf{\hat{n}}_{\left[0\right]}$ as projected onto the $xz$-plane. (Credit: Kim & Trippe (2021), reproduced with modifications.)
@@
In our pulsar emission model, a light ray of curvature radiation is emitted from the dipole magnetic field:
@@
$\mathbf{\hat{n}}_{\left[0\right]}=\beta\mathbf{\hat{B}}+\frac{\mathbf{\Omega}\times\mathbf{r}}{c},$ (13)
@@
$\hat{n}_{y\left[0\right]}\approx\frac{\Omega}{c}\left[\frac{\sin\alpha s}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}+x_{\mathrm{o}}\right]+\mathcal{O}\left(\phi^{2},\left(\Omega r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega r_{\mathrm{o}}/c\right)\right),$ (17)
@@
$y\approx\int\hat{n}_{y\left[0\right]}\,\mathrm{d}s+y_{\mathrm{o}}=\frac{\Omega}{c}\left[\frac{\sin\alpha s^{2}}{2\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}+x_{\mathrm{o}}s\right]+y_{\mathrm{o}},$ (20)
@@
$\hat{n}_{y}=\hat{n}_{y\left[0\right]}\approx\frac{\Omega r_{\mathrm{o}}}{c}\left[\frac{\sin\alpha\lambda}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}+\sin\theta_{\mathrm{o}}\right],$ (31)
@@
$Y\equiv\int_{0}^{s}\hat{n}_{y}\,\mathrm{d}s=r_{\mathrm{o}}\int_{0}^{\lambda}\hat{n}_{y\left[0\right]}\,\mathrm{d}\lambda\approx\frac{\Omega r_{\mathrm{o}}^{2}}{c}\left[\frac{\sin\alpha\lambda^{2}}{2\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}+\sin\theta_{\mathrm{o}}\lambda\right],$ (34)
@@
$\left.\hskip 33.0pt+8\cos\left(\theta_{\mathrm{o}}-\alpha\right)\left(141\cos^{4}\left(\theta_{\mathrm{o}}-\alpha\right)+86\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+13\right)_{\,}^{\,}\!\!\right]$
@@
For example, with $r_{\mathrm{o}}=2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\Omega=2\pi\times 10^{2}\,\mathrm{Hz}$ and $\eta_{2}B_{\mathrm{o}}^{2}\approx 4.29\times 10^{-5}$, we plot a trajectory of our light ray $\left(X/r_{\mathrm{o}},Y/r_{\mathrm{o}},Z/r_{\mathrm{o}}\right)$ for $0\leq\lambda\leq 10$ on a logarithmic scale, as shown in Fig. 2a. Also, for intuitive visualization, in Fig. 2b is plotted the trajectory on a linear scale, with the quantum refraction effect fairly exaggerated by $\eta_{2}B_{\mathrm{o}}^{2}\sim 10^{4}$, which is $10^{9}$ times as large as an actual order $\sim 10^{-5}$. Note, in particular, that the trajectory is deflected from a straight line as viewed in the $xz$-plane (due to the quantum refraction effect), and at the same time that it follows a parabolic path in another plane perpendicular to the $xz$-plane (due to the rotational effect of the pulsar magnetosphere); therefore, the light ray follows a three-dimensional twisted curve. In Appendix A we provide a detailed discussion of the properties of this curve in reference to the Frenet–Serret formulas (Spivak (1999)).
Figure 2: (a) A trajectory of the light ray $\left(X/r_{\mathrm{o}},Y/r_{\mathrm{\ o}},Z/r_{\mathrm{o}}\right)$ plotted against $0\leq\lambda\leq 10$ on a logarithmic scale; the red solid curve and the blue dashed curve represent the total trajectory (classical trajectory $+$ quantum correction) and the classical trajectory, respectively, (b) The trajectory plotted on a linear scale for intuitive visualization, with the quantum refraction effect fairly exaggerated by $\eta_{2}B_{\mathrm{o}}^{2}\sim 10^{4}$, which is $10^{9}$ times as large as an actual order $\sim 10^{-5}$.